OpenStudy (anonymous):

What is the directrix of x^2-28y?

jimthompson5910 (jim_thompson5910):

where is the equal sign?

OpenStudy (anonymous):

Let me check again

OpenStudy (anonymous):

Oh oops the minus is supposed to be an equal sign.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

so it's x^2 = 28y ?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

x^2 = 28y (x-0)^2 = 4*7(y-0) which is in the form (x-h)^2 = 4p(y-k) so the focal distance is p = 7 units

jimthompson5910 (jim_thompson5910):

the vertex is (0,0), so this means that the directrix is y = -7 since you go down 7 units you go down 7 units because this is a parabola that opens upward

OpenStudy (anonymous):

Thanks! :)

jimthompson5910 (jim_thompson5910):

yw